# Low-Rank Tensor Decomposition over Finite Fields

### v3

@@ 1 Introduction
This equation is also written as $T=\sum_{r=0}^{R-1}A_{r,:}\times B_{r,:}\times C_{r,:}$, where $M_{r,:}$ denotes the slice $[M_{r,j}]_{j}$ and $\times$ denotes the outer product: for arbitrary tensors $X$ and $Y$, $(X\times Y)_{i_{0},\dots i_{k-1},j_{0},\dots j_{\ell-1}}:=X_{i_{0},\dots i_{k-1}}Y_{j_{0},\dots j_{\ell-1}}$.
@@ 1.1 Results
For $R\leq 4$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(f(|\mathbb{F}|,R)n^{3})$ time, for some function $f$.
We show tighter runtime analysis in Section 2.3.1.
@@ 2.1 Preliminary results about matrix factorizations
Using the well-known fact that row-rank always equals column-rank, we can combine Lemma 4 with its transpose to get a stronger condition:
@@
For $U\in\mathbb{F}^{m\times r}$ and $V\in\mathbb{F}^{r\times n}$, let $r^{\prime}:=\mathrm{rank}\left(UV\right)$. Then $\left(\mathrm{rank}\left(U\right)\leq r^{\prime}\ \lor\ \mathrm{rank}\left(V\right)<r\right)\ \land\ \left(\mathrm{rank}\left(U\right)<r\ \lor\ \mathrm{rank}\left(V\right)\leq r^{\prime}\right)$.
@@ 2.2.4 Rows sharing two common columns
We can solve several easy cases on $\mathrm{rank}\left(D_{0}\right)$ and $\mathrm{rank}\left(D_{1}\right)$. By the rank precondition, we must have $\mathrm{rank}\left(D_{0}\right)\leq 3$ and $\mathrm{rank}\left(D_{1}\right)\leq 3$.
@@
    We can enumerate each possible rank-3 factorization of $D_{0}$, which fixes $(M_{0},M_{2},M_{3})$, then check if $\mathrm{rank}\left(D_{1}-M_{2}-zM_{3}\right)\leq 1$; if so, we have a solution.
@@
    Likewise, we can enumerate each rank-3 factorization of $D_{1}$, which fixes $(M_{1},M_{2},M_{3})$, then check if $\mathrm{rank}\left(D_{0}-M_{2}-zM_{3}\right)\leq 1$.
@@
      We can construct an arbitrary rank-2 factorization $D_{1}=x_{0}y_{0}+x_{1}y_{1}$, where $x_{i}$ are column-vectors and $y_{i}$ are row-vectors. We then try setting $M_{0}=D_{0},\ M_{1}=x_{0}y_{0},\ M_{2}+M_{3}=O,\ M_{2}+zM_{3}=x_{1}y_{1}$. This yields $M_{2}=-M_{3}$, so $(z-1)M_{3}=x_{1}y_{1}$. Since $z\neq 1$, $M_{3}=\frac{1}{z-1}x_{1}y_{1}$, so we again have a solution:
@@
      Likewise, we can factorize $D_{0}=x_{0}y_{0}+x_{1}y_{1}$, then attempt setting $M_{0}=x_{0}y_{0},\ M_{1}=D_{1}$ and imposing $M_{2}+M_{3}=x_{1}y_{1},\ M_{2}+zM_{3}=O$. We then have the solution $M_{3}=-\frac{1}{z-1}x_{1}y_{1},\ M_{2}=-zM_{3}$.
This leaves $(\mathrm{rank}\left(D_{0}\right),\mathrm{rank}\left(D_{1}\right))=(2,2)$ as the only remaining case. To solve it, we first analyze the matrices $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right],\ \Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$, obtained by concatenating $D_{0}$ and $D_{1}$ vertically and horizontally, respectively.
Recall that $M_{r}:=B_{r,:}\times C_{r,:}$ in our original problem; we will define column-vectors $u_{r}:=\begin{bmatrix}\vdots\\ B_{r,i}\\ \vdots\end{bmatrix}_{i}$ and row-vectors $v_{r}:=\begin{bmatrix}\cdots&C_{r,i}&\cdots\end{bmatrix}_{i}$ for brevity. Using the system of equations, $\Phi=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right]\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]$ and $\Theta=\left[\begin{array}[]{c|c|c|c}u_{0}&\overrightarrow{0}&u_{2}&u_{3}\\ \hline\cr\overrightarrow{0}&u_{1}&u_{2}&zu_{3}\end{array}\right]\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right]$, so a necessary condition for a solution to exist is that $\mathrm{rank}\left(\Phi\right)\leq 4$ and $\mathrm{rank}\left(\Theta\right)\leq 4$.
We can solve two easy cases of $\mathrm{rank}\left(\Phi\right)$:
  • If $\mathrm{rank}\left(\Phi\right)\leq 2$:
    We can find some arbitrary rank-2 factorization $\Phi=\left[\begin{array}[]{c|c}x_{0}&x_{1}\end{array}\right]\left[\begin{array}[]{c|c}y_{0}&z_{0}\\ \hline\cr y_{1}&z_{1}\end{array}\right]$ (padding with zeros if needed) for $x_{i}\in\mathbb{F}^{n\times 1}$ and $y_{i},\ z_{i}\in\mathbb{F}^{1\times n}$; this implies $D_{0}=x_{0}y_{0}+x_{1}y_{1},\ D_{1}=x_{0}z_{0}+x_{1}z_{1}$.
    To obtain a solution, we will attempt setting $M_{0}=x_{0}y_{0},\ M_{1}=x_{0}z_{0}$; then we need $M_{2},M_{3}$ to satisfy $M_{2}+M_{3}=x_{1}y_{1},\ M_{2}+zM_{3}=x_{1}z_{1}$. Since $x_{1}$ appears on the right-hand side of both of these equations, we will attempt further setting $u_{2}=u_{3}=x_{1}$. Then we have $M_{2}+M_{3}=x_{1}(v_{2}+v_{3}),\ M_{2}+zM_{3}=x_{1}(v_{2}+zv_{3})$, so we need $v_{2}+v_{3}=y_{1},\ v_{2}+zv_{3}=z_{1}$. Since $z\neq 1$, we can set $v_{3}=\frac{z_{1}-y_{1}}{z-1}$ and $v_{2}=y_{1}-v_{3}$. Thus, we can always find a solution, which written in full (in terms of $x_{i},y_{i},z_{i}$) is:
    $\left(M_{0},M_{1},M_{2},M_{3}\right)=\left(x_{0}y_{0},x_{0}z_{0},x_{1}\left(y_{1}-\frac{z_{1}-y_{1}}{z-1}\right),x_{1}\left(\frac{z_{1}-y_{1}}{z-1}\right)\right).$
  • Else if $\mathrm{rank}\left(\Phi\right)=4$:
    We can enumerate all possible rank-4 factorizations of $\Phi=U_{*}W_{*}$ and check if there exist $u_{i},v_{i}$ s.t. $\left(\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right],\ \left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]\right)=(U_{*},\ W_{*})$  [^6] It suffices to solve $\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$, which can be done in $O(n)$ time because each $v_{i}$ is already forced to be a specific submatrix of $W_{*}$..
By transposing the system, we can solve the cases $\mathrm{rank}\left(\Theta\right)\leq 2$ and $\mathrm{rank}\left(\Theta\right)=4$ in a similar manner.
Thus, the remaining case we must solve is:
If $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2$ and $\mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3$:
We have the following matrix factorizations:
$\Phi=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right]\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right],\ \Theta=\left[\begin{array}[]{c|c|c|c}u_{0}&\overrightarrow{0}&u_{2}&u_{3}\\ \hline\cr\overrightarrow{0}&u_{1}&u_{2}&zu_{3}\end{array}\right]\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right],$
$D_{0}=\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix},\ D_{1}=\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}.$
For brevity, we define $U:=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right],\ W:=\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]$.
Using Corollary 1 and the property that $\mathrm{rank}\left(AB\right)\leq\mathrm{min}\left(\mathrm{rank}\left(A\right),\mathrm{rank}\left(B\right)\right)$ for any matrices $A$ and $B$, we have:
  • $\mathrm{rank}\left(\Phi\right)=3\Rightarrow(\mathrm{rank}\left(U\right),\mathrm{rank}\left(W\right))\in\left\{(3,4),(4,3)\right\}$
  • $\mathrm{rank}\left(D_{0}\right)=2\Rightarrow\left(\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\right),\mathrm{rank}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)\right)\in\left\{(2,3),(3,2)\right\}$
  • $\mathrm{rank}\left(D_{0}\right)=2\Rightarrow\left(\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\right),\mathrm{rank}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}\right)\right)\in\left\{(2,3),(3,2)\right\}$
    • Since $z\neq 0$, we can replace $\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{bmatrix}$ with $\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}$ in the above statement.
We then split casework on $\mathrm{rank}\left(U\right)$:
  • If $\mathrm{rank}\left(U\right)=3$:
    By rank factorization, $\exists U^{\prime}\in\mathbb{F}^{n\times 3},\ F\in\mathbb{F}^{3\times 4}$ s.t. $U=U^{\prime}F$.
@@
    Then $\Phi=UW=U^{\prime}(FW)$ forms a rank-3 factorization of a rank-3 matrix. By Lemma 3, if we can choose some arbitrary $U_{*}\in\mathbb{F}^{n\times 3},\ W_{*}\in\mathbb{F}^{3\times 2n}$ s.t. $\Phi=U_{*}W_{*}$ (e.g., $W_{*}=\mathrm{rref}\left(\Phi\right)$), we are guaranteed that $\exists X\in\mathrm{GL}\left(3,\ \mathbb{F}\right)$ s.t. $(U^{\prime},FW)=(U_{*}X^{-1},XW_{*})$.
    Rearranging terms, we have $U=U_{*}X^{-1}F$ and $X^{-1}FW=W_{*}$. Defining $F^{\prime}:=X^{-1}F$ then yields $U=U_{*}F^{\prime},\ F^{\prime}W=W_{*}$. Since $U_{*},\ W_{*}$ are already known (as they are determined by $\Phi$) and $F^{\prime}\in\mathbb{F}^{3\times 4}$, we can enumerate all possible $F^{\prime}$ and check if $F^{\prime}W=W_{*}$ is solvable for the row-vectors $v_{i}$, which forms a linear system.
  • Else: $\mathrm{rank}\left(U\right)=4$
    The columns of $U$ are linearly independent, so $\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]\right)=3$ and $\mathrm{rank}\left(\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]\right)=3$. This implies $\mathrm{rank}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=2$ and $\mathrm{rank}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=2$.
    We also have $\mathrm{rank}\left(W\right)=3$, so $\exists\begin{bmatrix}\alpha&\beta&\gamma&\delta\end{bmatrix}\neq\overrightarrow{0}$ s.t. $\begin{bmatrix}\alpha&\beta&\gamma&\delta\end{bmatrix}W=\overrightarrow{0}$, and thus:
    $\left\{\begin{array}[]{cccc}\alpha v_{0}&&+\gamma v_{2}+\delta v_{3}&=\overrightarrow{0}\\ &\beta v_{1}&+\gamma v_{2}+z\delta v_{3}&=\overrightarrow{0}\end{array}\right\}.$
    If $\alpha\neq 0$, then $v_{0}\in\mathrm{rowspan}\left(\begin{bmatrix}v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$Rightarrow $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=\mathrm{rowspan}\left(\begin{bmatrix}v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$. If $\beta\neq 0$, then $v_{1}\in\mathrm{rowspan}\left(\begin{bmatrix}v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$Rightarrow $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)=\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$. Either case implies that $\mathrm{rowspan}\left(\begin{bmatrix}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{bmatrix}\right)$ is equal to the row-span of a rank-2 matrix, so $\mathrm{rank}\left(\Theta\right)\leq 2$: contradiction.
    Thus, we need $\alpha=0$ and $\beta=0$, implying $(\gamma,\delta)\neq(0,0)$ and $\left\{\begin{array}[]{cc}\gamma v_{2}+\delta v_{3}&=\overrightarrow{0}\\ \gamma v_{2}+z\delta v_{3}&=\overrightarrow{0}\end{array}\right\}$. Subtracting between the two equations yields $(z-1)\delta v_{3}=\overrightarrow{0}$. Since $z\neq 1$, $\delta v_{3}=\overrightarrow{0}$, so $\gamma v_{2}=\overrightarrow{0}$.
    Since $\gamma$ and $\delta$ are not both zero, at least one of $v_{2}$ and $v_{3}$ must be $\overrightarrow{0}$, so at least one of $M_{2}$ and $M_{3}$ is $O$. Either case means that the system reverts to the rows-sharing-one-common column case.
Summary of Casework
Below we summarize our full casework for solving the system $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{2}+M_{3}\\ D_{1}=&&M_{1}&+M_{2}+zM_{3}\end{array}\right\}$, where $z\neq 0,1$. We omit computation details for brevity, as they have already been mentioned above. As a reminder, we define $\Phi:=\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right],\ \Theta:=\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]$.
  • If $\mathrm{rank}\left(\Phi\right)>4$ or $\mathrm{rank}\left(\Theta\right)>4$ or $\mathrm{rank}\left(D_{0}\right)>3$ or $\mathrm{rank}\left(D_{1}\right)>3$:
    • No solution
  • Else if $\mathrm{rank}\left(D_{0}\right)=3$ or $\mathrm{rank}\left(D_{1}\right)=3$:
    • Enumerate rank factorizations of $D_{0}$ (or $D_{1}$)
  • Else if $\mathrm{rank}\left(D_{0}\right)\leq 1$ or $\mathrm{rank}\left(D_{1}\right)\leq 1$:
    • Set $M_{0}=D_{0}$ (or $M_{1}=D_{1}$) and use an arbitrary rank-2 factorization of $D_{1}$ (or $D_{0}$)
  • Else if $\mathrm{rank}\left(\Phi\right)=4$ or $\mathrm{rank}\left(\Theta\right)=4$:
    • Enumerate rank factorizations of $\Phi$ (or $\Theta$)
  • Else if $\mathrm{rank}\left(\Phi\right)\leq 2$ or $\mathrm{rank}\left(\Theta\right)\leq 2$:
    • Use an arbitrary rank-2 factorization of $\Phi$ (or $\Theta$)
@@
  • Else: we must have $\mathrm{rank}\left(\Phi\right)=\mathrm{rank}\left(\Theta\right)=3$ and $\mathrm{rank}\left(D_{0}\right)=\mathrm{rank}\left(D_{1}\right)=2$:
    • Find an arbitrary rank-3 factorization $\Phi=U_{*}W_{*}$, then solve $F\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]=W_{*}$ for $F\in\mathbb{F}^{3\times 4}$ and $v_{i}\in\mathbb{F}^{1\times n}$
    • if no solution is found: solve $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{2}\\ D_{1}=&&M_{1}&+M_{2}\end{array}\right\}$
    • if no solution is found: solve $\left\{\begin{array}[]{cccc}D_{0}=&M_{0}&&+M_{3}\\ D_{1}=&&M_{1}&+zM_{3}\end{array}\right\}$

### v4

@@ 1 Introduction
This equation is also written as $T=\sum_{r=0}^{R-1}A_{r,:}\times B_{r,:}\times C_{r,:}$, where $M_{r,:}$ denotes the $r$-th row of $M$ and $\times$ denotes the outer product: for arbitrary tensors $X$ and $Y$, $(X\times Y)_{i_{0},\dots i_{k-1},j_{0},\dots j_{\ell-1}}:=X_{i_{0},\dots i_{k-1}}Y_{j_{0},\dots j_{\ell-1}}$.
@@ 1.1 Results
For fixed $R\leq 4$, finding a rank-$R$ decomposition over a finite field $\mathbb{F}$ of an $n\times n\times n$ tensor, or determining that it does not exist, can be done in $O(n^{3}+f(|\mathbb{F}|,R)n^{2})$ time, for some function $f$.
@@ 2.1 Preliminary results about matrix factorizations
Using the well-known properties $\mathrm{rank}\left(A\right)=\mathrm{rank}\left(A^{\intercal}\right)$ and $\mathrm{rank}\left(AB\right)\leq\mathrm{min}\left(\mathrm{rank}\left(A\right),\mathrm{rank}\left(B\right)\right)$ for arbitrary matrices $A$ and $B$, we can combine Lemma 4 with its transpose to get a stronger condition:
@@
For $U\in\mathbb{F}^{m\times r}$ and $V\in\mathbb{F}^{r\times n}$, let $r^{\prime}:=\mathrm{rank}\left(UV\right)$. Then $\left(\mathrm{rank}\left(U\right)\leq r^{\prime}\ \lor\ \mathrm{rank}\left(V\right)<r\right)\ \land\ \left(\mathrm{rank}\left(U\right)<r\ \lor\ \mathrm{rank}\left(V\right)\leq r^{\prime}\right)\ \land\ \mathrm{rank}\left(U\right)\geq r^{\prime}\ \land\ \mathrm{rank}\left(V\right)\geq r^{\prime}$.
@@ 2.2.4 Rows sharing two common columns
We can solve several easy cases on $\mathrm{rank}\left(D_{0}\right)$ and $\mathrm{rank}\left(D_{1}\right)$. WLOG assume the rank precondition: $\mathrm{rank}\left(D_{0}\right)\leq 3\ \land\ \mathrm{rank}\left(D_{1}\right)\leq 3$.
@@
    Enumerate each of the $O(|\mathbb{F}|^{9})$ possible full-rank factorizations of $D_{0}$, which fixes $(M_{0},M_{2},M_{3})$;
    then check if $\mathrm{rank}\left(D_{1}-M_{2}-zM_{3}\right)\leq 1$.
@@
    Likewise, enumerate each full-rank factorization of $D_{1}$, which fixes $(M_{1},M_{2},M_{3})$;
    then check if $\mathrm{rank}\left(D_{0}-M_{2}-zM_{3}\right)\leq 1$.
@@
      Set $M_{0}=D_{0}$;
      construct an arbitrary rank-2 factorization $D_{1}=x_{0}y_{0}+x_{1}y_{1}$, where $x_{i}$ are column-vectors and $y_{i}$ are row-vectors; then set $M_{1}=x_{0}y_{0}$.
      We now must solve $M_{2}+M_{3}=O,\ M_{2}+zM_{3}=x_{1}y_{1}$. This yields $M_{2}=-M_{3}$, and since $z\neq 1$, $M_{3}=\frac{1}{z-1}x_{1}y_{1}$, so we always have a solution:
@@
      Likewise, set $M_{1}=D_{1}$, factorize $D_{0}=x_{0}y_{0}+x_{1}y_{1}$, set $M_{0}=x_{0}y_{0}$. Then we must solve $M_{2}+M_{3}=x_{1}y_{1},\ M_{2}+zM_{3}=O$, which can be done by setting $M_{3}=-\frac{1}{z-1}x_{1}y_{1},\ M_{2}=-zM_{3}$.
This leaves $\mathrm{rank}\left(D_{0}\right)=2\ \land\ \mathrm{rank}\left(D_{1}\right)=2$ as the only remaining case. We can substitute $M_{i}=u_{i}v_{i}$, with $u_{i}\in\mathbb{F}^{n\times 1},\ v_{i}\in\mathbb{F}^{1\times n}$, since $\mathrm{rank}\left(M_{i}\right)\leq 1$. Then solving the system is equivalent to solving $D_{0}=\underbrace{\left[\begin{array}[]{c|c|c}u_{0}&u_{2}&u_{3}\end{array}\right]}_{=:U_{0}}\underbrace{\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right]}_{=:V_{0}}$ and $D_{1}=\underbrace{\left[\begin{array}[]{c|c|c}u_{1}&u_{2}&u_{3}\end{array}\right]}_{=:U_{1}}\underbrace{\left[\begin{array}[]{c}v_{1}\\ \hline\cr v_{2}\\ \hline\cr zv_{3}\end{array}\right]}_{=:V_{1}}$ simultaneously. Invoking Corollary 1, we have:
  • $\mathrm{rank}\left(D_{0}\right)=2\Rightarrow(\mathrm{rank}\left(U_{0}\right),\mathrm{rank}\left(V_{0}\right))\in\left\{(3,2),(2,3),(2,2)\right\}$
  • $\mathrm{rank}\left(D_{1}\right)=2\Rightarrow(\mathrm{rank}\left(U_{1}\right),\mathrm{rank}\left(V_{1}\right))\in\left\{(3,2),(2,3),(2,2)\right\}$
Collectively, we have $\left(\mathrm{rank}\left(U_{0}\right)=2\ \lor\ \mathrm{rank}\left(V_{0}\right)=2\right)\ \land\ \left(\mathrm{rank}\left(U_{1}\right)=2\ \lor\ \mathrm{rank}\left(V_{1}\right)=2\right)$.
Although the factorizations $D_{i}=U_{i}V_{i}$ are not full-rank, at least one of $U_{i}$ and $V_{i}$ must have the same rank as $D_{i}$, which motivates the following lemma:
Lemma 5.
@@
For $U\in\mathbb{F}^{m\times r},\ V\in\mathbb{F}^{r\times n}$, define $D:=UV$ and $r^{\prime}:=\mathrm{rank}\left(D\right)$. Let $D=U_{*}V_{*}$ be a full-rank factorization of $D$, with $U_{*}\in\mathbb{F}^{m\times r^{\prime}},\ V_{*}\in\mathbb{F}^{r^{\prime}\times n}$. Suppose furthermore that $\mathrm{rank}\left(U\right)=r^{\prime}$. Then $U=U_{*}G$ for some $G\in\mathbb{F}^{r\times r^{\prime}}$.
Proof.
By Lemma 3, there exists a full-rank factorization of $U=U^{\prime}F$, where $U^{\prime}\in\mathbb{F}^{m\times r^{\prime}},\ F\in\mathbb{F}^{r^{\prime}\times r}$. Then $D=U^{\prime}(FV)$ is a full-rank factorization of $D$, so by Lemma 3 again, $\exists X\in\mathrm{GL}\left(r^{\prime},\ \mathbb{F}\right)$ s.t. $(U^{\prime},FV)=(U_{*}X,X^{-1}V_{*})$, which yields $U=U_{*}(XF)$. 666 Notice that $V$ is much less restricted, as it only has to satisfy is $(XF)V=V_{*}$, which could have exponentially many solutions w.r.t. $m,n$.
Thus, we have if $UV$ is a rank-$r$ factorization of a rank-$r^{\prime}$ matrix, and that $\mathrm{rank}\left(U\right)=r$, then there are only $O(|\mathbb{F}|^{rr^{\prime}})$ possibilities for $U$, and each can be enumerated in polynomial time via matrix multiplication. By matrix transpose, a similar result applies to $V$ if $\mathrm{rank}\left(V\right)=r$. Once again, we avoid exponential dependence on the dimensions of the matrices.
We can now solve our last case by splitting it into the following four subcases:
  • If $\mathrm{rank}\left(U_{0}\right)=2\ \land\ \mathrm{rank}\left(U_{1}\right)=2$:
    Enumerate all possible pairs $(U_{0},U_{1})$, which fixes the column-vectors $u_{i}$;
    for each pair, if it does not cause conflicts on what each $u_{i}$ gets fixed to (i.e. if $(U_{1})_{:,2}=(U_{0})_{:,2}$ and $(U_{1})_{:,3}=(U_{0})_{:,3}$), solve the linear system $\left[\begin{array}[]{c}D_{0}\\ \hline\cr D_{1}\end{array}\right]=\left[\begin{array}[]{c|c|c|c}u_{0}&\overrightarrow{0}&u_{2}&u_{3}\\ \hline\cr\overrightarrow{0}&u_{1}&u_{2}&zu_{3}\end{array}\right]\left[\begin{array}[]{c}v_{0}\\ \hline\cr v_{1}\\ \hline\cr v_{2}\\ \hline\cr v_{3}\end{array}\right]$ for $v_{i}$.
  • Else if $\mathrm{rank}\left(V_{0}\right)=2\ \land\ \mathrm{rank}\left(V_{1}\right)=2$:
    Enumerate all possible $(V_{0},V_{1})$, which fixes the row-vectors $v_{i}$;
    • Notice that this step relies on $z$ being nonzero, as otherwise it is possible for $v_{3}$ to still have exponentially many possibilities.
    for each pair, if it does not cause conflicts on $v_{i}$, solve $\left[\begin{array}[]{c|c}D_{0}&D_{1}\end{array}\right]=\left[\begin{array}[]{c|c|c|c}u_{0}&u_{1}&u_{2}&u_{3}\end{array}\right]\left[\begin{array}[]{c|c}v_{0}&\overrightarrow{0}\\ \hline\cr\overrightarrow{0}&v_{1}\\ \hline\cr v_{2}&v_{2}\\ \hline\cr v_{3}&zv_{3}\end{array}\right]$ for $u_{i}$.
  • Else if $\mathrm{rank}\left(U_{0}\right)=2\ \land\mathrm{rank}\left(V_{1}\right)=2$:
    Enumerate all possible $(U_{0},V_{1})$, which fixes all variables except $u_{1}$ and $v_{0}$;
    for each pair, solve $u_{0}v_{0}=D_{0}-u_{2}v_{2}-u_{3}v_{3}$ for $v_{0}$ and $u_{1}v_{1}=D_{1}-u_{2}v_{2}-zu_{3}v_{3}$ for $u_{1}$.
@@
  • Else: must have $\mathrm{rank}\left(U_{1}\right)=2\ \land\mathrm{rank}\left(V_{0}\right)=2$:
    Enumerate all possible $(U_{1},V_{0})$;
    for each pair, solve $u_{0}v_{0}=D_{0}-u_{2}v_{2}-u_{3}v_{3}$ for $u_{0}$ and $u_{1}v_{1}=D_{1}-u_{2}v_{2}-zu_{3}v_{3}$ for $v_{1}$.
In each subcase, when enumerating matrix pairs, we apply Lemma 5 with $r=3$ and $r^{\prime}=2$, so each matrix in the pair has $O(|\mathbb{F}|^{6})$ many possibilities, resulting in the pair having $O(|\mathbb{F}|^{12})$ many possibilities.